# Berry monopoles and divergent quantum metric in systems with non-integer band dispersion

### v1

@@ III Berry curvature
This integration is performed over the entire $\mathbf{k}$-space since our model, as described by Eq. (1), is based on the continuum Dirac Hamiltonian. It is important to note that the Dirac Hamiltonian does not possess a proper Brillouin zone; instead, its energy bands extend indefinitely [34, 35]. As a result, the quantity defined in Eq. (5) is not quantized to integer values and technically does not represent a Chern number, which is only defined for compact $\mathbf{k}$ spaces. However, the change in $C$ across a possible topological transition, as the gap varies from $m=-1$ to $m=1$, yields an integral value [34, 35] and can be interpreted as a change in the Chern number across this transition.
@@
$\Delta C=C|_{m=+1}-C|_{m=-1}~{}.$ (6)
As shown in Fig. 2 (a), we find that $\Delta C$ is quantized to the integer value $-1$, even for non-integer values of $\alpha>0.5$. In the interval $0.50<\alpha<$ 0.5, however, the Berry curvature spreads along the axes, requiring a larger momentum region for integration to accurately capture the change in the Chern number. This explains the observed deviation from $\Delta C=-1$ in the figure.
In the special case where $\alpha=0$, the situation is different. As seen in Fig. 1 (e), the Berry curvature is uniformly zero across momentum space, remaining so even as $m$ varies from negative to positive. This results in $\Delta C=0$, indicating that there is no topological transition when $m$ is varied across zero for $\alpha=0$. This conclusion is consistent with our previous discussion, where we noted the absence of band gap closing at $m=0$ for $\alpha=0$.
IV Orbital magnetization
To explore the possible experimental consequences of the Berry curvature distribution in systems with non-integer energy-momentum dispersion, we calculate the orbital magnetization [22, 36, 37]. In magic-angle graphene, for example, the orbital magnetism can be imaged using a scanning superconducting quantum interference device [38]. Here we have two contributions: the local orbital magnetization
$\mathbf{M}_{orb}^{loc}=\frac{1}{(2\pi)^{3}}\mathcal{I}m\sum_{n}\int_{E_{n\mathbf{k}}\leq\mu}d^{2}k\langle\partial_{\mathbf{k}}\psi_{n,\mathbf{k}}|\times H_{\mathbf{k}}|\partial_{\mathbf{k}}\psi_{n,\mathbf{k}}\rangle$ (7)
@@ IV Orbital magnetization
$\mathbf{M}_{orb}^{it}=$ $\frac{1}{(2\pi)^{3}}\mathcal{I}m\sum_{n}\int_{E_{n\mathbf{k}}\leq\mu}d^{2}k\langle\partial_{\mathbf{k}}\psi_{n,\mathbf{k}}|\times(E_{n,\mathbf{k}}$
$-2\mu)|\partial_{\mathbf{k}}\psi_{n,\mathbf{k}}\rangle~{},$ (8)
where $\mu$ is the chemical potential.
@@
Here, we consider the conduction band only. The total orbital magnetization $M_{orb}$ has an explicit proportionality with the integral of the Berry curvature [39]
$\mathbf{M}_{orb}=\frac{2}{(2\pi)^{2}}\int_{E_{\mathbf{k}}\leq\mu}d^{2}k\mu\mathbf{\Omega}(\mathbf{k})~{}.$ (9)
@@
We show our results on Fig. 2 (b). The orbital magnetization is constant in the range $\alpha\gtrsim 0.4$. This is because while the Berry curvature is redistributed inside the momentum region $E_{\mathbf{k}}\leq\mu$ as $\alpha$ is varied, it does not spread outside the integration region, rendering the integral constant. We found that this constant magnetization persists for larger $\alpha$, but the numerical integration must be refined as the Berry curvature becomes highly localized into four monopoles. For $0.40<\alpha<$ 0.4, some of the Berry curvature leaks out of the integration region as shown in Fig. 1 (f), resulting in decreased orbital magnetization magnitude. At exactly $\alpha=0$, the orbital magnetization contributions are zero since the Berry curvature vanishes identically in the entire momentum space as previously shown in Fig. 1 (e). This behavior holds even if $m$ is varied from negative to positive, which is consistent with our result above that there is no topological transition for $\alpha=0$.
V Quantum metric
Figure 3: The plots illustrate the trace (a)-(d) and the determinant (e)-(h) of the quantum metric for the conduction band across different values of $\alpha$, with the band gap parameter set to $m=0.5$. In panels (b) and (f), we used $\alpha=0.3$, while in panels (d) and (h), $\alpha=2.5$.
We now calculate the quantum metric
$g_{ij}(\mathbf{k},\alpha)_{\pm}=$ $\langle\partial_{i}\psi_{\pm}(\mathbf{k},\alpha)|\partial_{j}\psi_{\pm}(\mathbf{k},\alpha)\rangle$
$-A_{i}(\mathbf{k},\alpha)_{\pm}A_{j}(\mathbf{k},\alpha)_{\pm}~{},$ (10)
@@ V Quantum metric
where $\mathbf{A}(\mathbf{k},\alpha)_{\pm}$ is the Berry connection and the $\pm$ index labels the upper and lower band cases. This gives us the information about the overlap of the states in $\mathbf{k}$ space. We show on Fig. 3 the determinant and trace of the quantum metric. It is sufficient to show the metric invariants for the upper band since the lower band have identical quantum metric components. For the flat band, $\alpha=0$, the quantum metric is zero everywhere in the momentum space as shown in Fig. 3 (a) and (e). This tells us that all the states have zero overlap and are therefore well-localized. We found similar results when $m$ is varied across $m=0$, which is consistent with previous results that there is no band gap closing, Fig. 1 (a), and that there is no change in the Chern number, Fig. 2 (b). Hence, for the flat band case, our model describes a trivial insulator for all values of $m$.
@@
For non-integer dispersions, $10<\alpha<$ 1, the quantum metric components are divergent along the momentum space axes as shown on Fig. 3 (b) and (f). This mirrors the non-analytic regions of the band dispersion and the Berry curvature, Figs. 1 (b) and (f), respectively. The states along the axes in this case are therefore highly overlapping.
For the massive Dirac band case shown in Figs. 3 (c) and (g), $\alpha=1$, we found the typical isotropic and smooth peak at the center of the momentum space. For $\alpha>1$, Figs. 3 (d) and (h), we found four smooth peaks, one for each quadrant in the momentum space. These peaks are found at the same locations of the four “petals”of the Berry curvature in Fig. 1 (h), which, in turn, as discussed above, are located at the corners of the rectangular-like surface of the energy band. In all cases, we found that the peaks in the quantum metric occurs at the locations where the energy band curvature is high, implying that the states in these regions are highly overlapping.
VI Conclusions
@@
To summarize, our calculations reveal that a non-integer dispersion introduces significant and novel effects on band geometry and topology. Specifically, we observe that the Berry curvature and quantum metric become concentrated and can even diverge in momentum space regions where the energy band exhibits high curvature or non-analytic behavior, such as the kinks along the axes for $10<\alpha<$ 1 and the corners for $\alpha>1$.
The divergence of the quantum metric along the axes for $10<\alpha<$ 1 is particularly noteworthy. In previously studied Hermitian systems, the quantum metric has primarily been considered in terms of corrections, enhancing various physical properties. For instance, recent work on the impact of isolated singularities in the quantum metric demonstrated an increase in the BKT temperature in flat-band superconductors [40]. While divergent quantum metrics have been predicted and observed in non-Hermitian systems, where they significantly influence wave packet dynamics [41, 42], our findings indicate that such divergences can also occur in Hermitian systems.
Moreover, while it is well understood that Berry monopoles form at band touchings, our results show that they can also arise at sharp band corners, even without band touchings. These sharp corners are specifically a consequence of non-integer dispersion. The distinct patterns of peaks and divergences in the quantum metric across different values of $\alpha$ suggest that state overlaps are markedly influenced by non-integer dispersion. This insight opens avenues for engineering desirable physical properties related to wave dynamics and transport in future research.

### v2

@@ III.2 Chern number
This integration is performed over the entire $\mathbf{k}$-space since our model, as described by Eq. (5), is based on the continuum Dirac Hamiltonian [37, 42]. As a result, the quantity defined in Eq. (27) is not quantized to integer values and technically does not represent a Chern number, which is only defined for compact $\mathbf{k}$ spaces. However, the change in $C$ across a possible topological transition, as the gap varies from $m=-1$ to $m=1$, can still be interpreted as a change in the Chern number across this transition [37, 42].
@@
$\Delta C=C|_{m=+1}-C|_{m=-1}.$ (28)
As shown in Fig. 3 (a), we find that $\Delta C$ is quantized to the integer value $-1$, even for non-integer values of $\alpha>0.5$. In the interval $0.50<\alpha<$ 0.5, however, the Berry curvature spreads along the axes, requiring a larger momentum region for integration to accurately capture the change in the Chern number. This explains the observed deviation from $\Delta C=-1$ in the figure. We investigate this change in Chern number further by deriving the closed-form formula for $C$.
For $\alpha=0$, Eq. (27) trivially vanishes since the integrand is zero, as can be seen from Eq. (19). We therefore focus on the case $\alpha\neq 0$. We obtain a closed-form expression by taking advantage of the C4-symmetry of the Berry curvature as established by Fig. 1. We get the total integral of $\Omega_{\alpha,xy}^{+}$ by quadrupling the contribution of the first quadrant. In this region, Eq. (21) simplifies to
$\Omega_{\alpha,xy}^{+}=-\frac{m\alpha^{2}(k_{x}k_{y})^{\alpha-1}}{2(k_{x}^{2\alpha}+k_{y}^{2\alpha}+m^{2})^{3/2}}.$ (29)
The expression for $C$ becomes
[…]
$=-\frac{1}{\pi}\tan^{-1}\Bigg{(}\dfrac{k_{y}^{\alpha}}{m}\Bigg{)}\Bigg{|}_{0}^{\infty}$ (32)
$=-\frac{\text{sgn}(m)}{2},$ (33)
which shows that $\alpha$ drops out in the final expression. This gives a quantized change in Chern number Eq. (28)
$\Delta C=-1,$ (34)
independent of $\alpha$, provided that $\alpha>0$. Note that this result is similar to that of the standard Dirac Hamiltonian $\alpha=1$ [37], thereby showing that the fractional dispersion does not change the topological property $\Delta C$.
For $\alpha=0$, the situation is different. As seen in Fig. 1 (e), the Berry curvature is uniformly zero across momentum space, remaining so even as $m$ varies from negative to positive. This can be seen in general from Eq. (19), where the Berry curvature is shown to be proportional to $\alpha^{2}$. This results in $\Delta C=0$, indicating that there is no topological transition when $m$ is varied across zero for $\alpha=0$. This is consistent with our previous discussion, where we noted the absence of band gap closing at $m=0$ for $\alpha=0$.
III.3 Orbital magnetization
Figure 4: The plots illustrate the trace (a)-(d) and the determinant (e)-(h) of the quantum metric for the conduction band across different values of $\alpha$, with the band gap parameter set to $m=0.5$. In panels (b) and (f), we used $\alpha=0.3$, with the domain of the plots zoomed in and the range displayed on a logarithmic scale to clearly highlight the divergence along the momentum space axes. In panels (d) and (h), we used $\alpha=2.5$.
To explore the physical consequence of the Berry curvature redistribution, we calculate the orbital magnetization [22, 43, 44], a quantity that is accessible in experiments. In magic-angle graphene, for example, the orbital magnetism can be imaged using a scanning superconducting quantum interference device [45]. We have two contributions: the local orbital magnetization
$\mathbf{M}_{orb}^{loc}=\frac{1}{(2\pi)^{2}}\mathcal{I}m\sum_{n}\int_{E_{n\mathbf{k}}\leq\mu}d^{2}k\langle\partial_{\mathbf{k}}\psi_{n,\mathbf{k}}|\times H_{\mathbf{k}}|\partial_{\mathbf{k}}\psi_{n,\mathbf{k}}\rangle$ (35)
@@ III.3 Orbital magnetization
$\mathbf{M}_{orb}^{it}=$ $\frac{1}{(2\pi)^{2}}\mathcal{I}m\sum_{n}\int_{E_{n\mathbf{k}}\leq\mu}d^{2}k\langle\partial_{\mathbf{k}}\psi_{n,\mathbf{k}}|\times(E_{n,\mathbf{k}}$
$-2\mu)|\partial_{\mathbf{k}}\psi_{n,\mathbf{k}}\rangle,$ (36)
@@
where $\mu$ is the chemical potential. Here, we consider the conduction band only. In conduction bands, the total orbital magnetization $M_{orb}$ was shown to be proportional to the integral of the Berry curvature [38].
@@
We show our results on Fig. 3 (b). The orbital magnetization is constant in the range $\alpha\gtrsim 0.4$. This is because while the Berry curvature is redistributed inside the momentum region $E_{\mathbf{k}}\leq\mu$ as $\alpha$ is varied, it does not spread outside the integration region, rendering the integral constant. We found that this constant magnetization persists for larger $\alpha$, but the numerical integration must be refined as the Berry curvature becomes highly localized into four monopoles. For $0.40<\alpha<$ 0.4, some of the Berry curvature leaks out of the integration region as shown in Fig. 1 (f), resulting in decreased orbital magnetization magnitude. At exactly $\alpha=0$, the orbital magnetization contributions are zero since the Berry curvature vanishes identically in the entire momentum space as previously shown in Fig. 1 (e).
It is interesting to compare our model and magnetization results with those of chirally stacked graphene $N$-layer systems, which exhibit a large exponent $N$ dispersion. Furthermore, as mentioned in the introduction, the application of an external electric field perpendicular to the layers can interpolate the dispersion between two integer values [8], potentially enabling the realization of fractional dispersion in multi-layer graphene systems.
In contrast to our finding that the magnetization is independent of $\alpha$, the magnetization in chirally stacked graphene is proportional to the exponent $N$ [46]. This discrepancy originates from the difference in the underlying Bloch Hamiltonians. The Hamiltonian for chirally stacked graphene is given by
$H_{N}=\frac{(v_{0}p)^{N}}{(-\gamma_{1})^{N-1}}[\cos(N\phi_{\mathbf{p}})\sigma_{x}+\sin(N\phi_{\mathbf{p}})\sigma_{y}]+\lambda\sigma_{z},$ (37)
where $\cos\phi_{\mathbf{p}}=\tau_{x}p_{x}/p$ and $\sin\phi_{\mathbf{p}}=p_{y}/p$.
In contrasts to our simpler 2 $\times$ 2 Hamiltonian model, Eq. (37) includes the layer pseudospin $\vec{\sigma}$, valley $\vec{\tau}$, as well as real spin $\vec{s}$, which gives rise to three distinct symmetry-broken ground states by replacing the term $\lambda\sigma_{z}$ with $\lambda\tau_{z}\sigma_{z}$, $\lambda s_{z}\sigma_{z}$, or $\lambda\tau_{z}s_{z}\sigma_{z}$.
For $N=1$, the model (37) reduces to two copies of Dirac Hamiltonian which is similar to our model when $\alpha=1$. The difference becomes more pronounced for $N=2$, where Eq. (37) becomes
$H=\frac{v_{0}^{2}}{(-\gamma_{1})}[(p_{x}^{2}-p_{y}^{2})\sigma_{x}+2p_{x}p_{y}\tau_{z}\otimes\sigma_{y}]+\lambda\sigma_{z}.$ (38)
It would be interesting to study the fractional generalization of the stacked graphene model:
$H_{\alpha}=\frac{(v_{0}p)^{\alpha}}{(-\gamma_{1})^{\alpha-1}}[\cos(\alpha\phi_{\mathbf{p}})\sigma_{x}+\sin(\alpha\phi_{\mathbf{p}})\sigma_{y}]+\lambda\sigma_{z},$ (39)
where $\alpha$ is not restricted to integer values, which we leave for future study.
IV Quantum metric
@@
We now examine the results of the quantum metric. This gives us the information about the overlap of the states in $\mathbf{k}$-space. We show on Fig. 4 the determinant and trace of the quantum metric which provide avenues for probing quantum geometry in experiment. The determinant supports an alternative method for characterizing topological invariants [47], while the trace has been demonstrated to measure the band drude weight through directional lattice oscillations [48, 49]. It is sufficient to show the metric invariants for the upper band since the lower band have identical quantum metric components as evident in Eq. (20). For the flat band, $\alpha=0$, the quantum metric is zero everywhere in the momentum space as shown in Fig. 4 (a) and (e). This tells us that all the states have complete overlap and are therefore identical. We found similar results when $m$ is varied across $m=0$, which is consistent with previous findings that there is no band gap closing, Fig. 1 (a), and that there is no change in the Chern number, Fig. 3 (b). Hence, for the flat band case, our model describes a trivial insulator for all values of $m$.
@@
For non-integer dispersions, $10<\alpha<$ 1, the quantum metric components are divergent along the momentum space axes as shown on Fig. 4 (b) and (f). This mirrors the non-analytic regions of the energy band dispersion and the Berry curvature, Figs. 1 (b) and (f), respectively. This can be understood by examining the exact forms for the trace and the determinant of the quantum metric,
$\text{Tr}(\mathbf{g}_{\alpha}^{\pm})=$ $\frac{\alpha^{2}}{4d(\mathbf{k},\alpha)^{4}}\bigg{[}|k_{x}|^{2(\alpha-1)}(|k_{y}|^{2\alpha}+m^{2})$
$+|k_{y}|^{2(\alpha-1)}(|k_{x}|^{2\alpha}+m^{2})\bigg{]}$ (40)
$\text{det}(\mathbf{g}_{\alpha}^{\pm})=$ $\frac{m^{2}\alpha^{4}}{16d(\mathbf{k},\alpha)^{8}}\bigg{[}m^{2}|k_{x}k_{y}|^{2(\alpha-1)}$
$+|kx|^{2(2\alpha-1)}|ky|^{2(\alpha-1)}$
$+|ky|^{2(2\alpha-1)}|kx|^{2(\alpha-1)}\bigg{]}.$ (41)
These expressions exhibit terms that are proportional to $|k_{i}|^{\alpha-1}$ which diverge along the momentum space axes $10<\alpha<$ 1. Just as in the case of the Berry curvature, the divergence of the quantum metric is also attributed to the divergent first derivatives over the axes of the fractional Bloch Hamiltonian for $10<\alpha<$ 1.
For $\alpha>1$, Figs. 4 (d) and (h), we found four smooth finite peaks, one for each quadrant in the momentum space. These peaks are found at the same locations of the four “petals”of the Berry curvature in Fig. 1 (h). The corresponding locations between the peaks of the quantum metric and the petals of the Berry curvature may be ascribed to the equalities,
$\text{Tr}(\mathbf{g^{\pm}})$ $=|\Omega_{xy}^{\pm}|$ (42)
$\sqrt{\text{det}(\mathbf{g^{\pm}})}$ $=\frac{|\Omega_{xy}^{\pm}|}{2},$ (43)
that are unique for two-band systems [50, 35]. Consistent with Eqs. (42) and (43), we found for all cases of $\alpha$ that the peaks in the quantum metric, just like that of the Berry curvature, occur at the locations where the energy band curvature is high.
For the massive Dirac band case shown in Figs. 4 (c) and (g), $\alpha=1$, we found the typical isotropic and smooth peak at the center of the momentum space.
While the transition from a finite metric for $\alpha>1$ to a divergent metric for $\alpha<1$ is found in the topological phase transitions in models like the Su-Schrieffer-Heeger model [51], we will argue that in our model, the regime $10<\alpha<$ 1 may be unphysical.
V Conclusions
@@
In summary, our calculations reveal that non-integer dispersion introduces significant and novel effects on quantum band geometry. Specifically, we observe that fractional dispersion redistributes both the Berry curvature and the quantum metric. These quantities become highly concentrated in momentum space regions where the energy band exhibits strong curvature, such as at the corners, for $\alpha>1$.
We now give our final comments for the region $10<\alpha<$ 1, reflecting from all the results that we have obtained. In this range, we observe that both the Berry curvature and the quantum metric diverge along the momentum axes. Additionally, the Fermi velocity, as described by Eqs. (11) and (12), as well as the momentum gradient of the Bloch eigenfunctions, also exhibit both divergences and discontinuities. The divergence in the latter implies that the center positions of the Wannier functions is located at infinity. These findings strongly indicate that the regime $10<\alpha<$ 1 may be unphysical. This conclusion is consistent with the path integral formulation of fractional quantum mechanics, where Brownian trajectories are replaced by Lévy flights. In this framework, the region $10<\alpha<$ 1 is excluded, as it leads to a divergent first moment [15]. Notably, even without invoking this framework of fractional quantum mechanics, the unphysical nature of this regime manifests as a discontinuity in the Fermi velocity and a divergence in the quantum geometric tensor.
It is well-established that Berry monopoles form at band touching points, with two key signatures. The first, and most significant, is the formation of topological charge in $\vec{d}$ space, which acts as a source or sink for the pseudo-magnetic field. The second is that the Berry curvature vanishes almost everywhere in the band, becoming concentrated at the band touching points. Our results reveal a separation of these two effects for fractional systems ($\alpha>1$). While the Berry monopole, corresponding to the topological charge at $\vec{d}=0$, remains located at the origin $\mathbf{k}=0$, the Berry curvature develops Dirac delta-like peaks at four distinct points, away from $\mathbf{k}=0$.
It is truly remarkable that quantum systems with fractional dispersion can be realized in a variety of media. In this work, we have demonstrated that the interplay between fractional quantum mechanics and quantum geometry leads to rich phenomena not observed in conventional systems with integer dispersion. The quantum geometric tensor, particularly its quantum metric component, remains largely unexplored even in traditional systems. This presents an exciting opportunity to investigate the physical implications of fractional dispersion on geometry. Future directions of research could include exploring the effects on Hall transport, superfluid weight in multi-band superconductors, and non-linear optical responses, among many other phenomena. We hope this work will inspire further investigations into the connections between fractional systems and geometry.